OpenStudy (anonymous):

compute the 9th derivative of f(x)=arctan(x^(3)/4) at x=0

OpenStudy (experimentx):

arc tan is a trouble as variable factory while taking n'th derivative.

OpenStudy (experimentx):

f1(x) = 1/(1+(x^3/4)^2) * 1/4 * 3x f2(x) = 3x/4( 1 /(1+(x^3/4)^2)^2*-1*3x^2/4 + 3/4(1+(x^3/4)^2) =

OpenStudy (experimentx):

looks it's a long way ... until we find a general pattern.

OpenStudy (anonymous):

I tried at first spending so much time listing them out and I could not find the pattern

OpenStudy (anonymous):

I think finding pattern and writing the equation is the most difficult part for me

OpenStudy (experimentx):

it's quite difficult .... it's going to be series.

OpenStudy (anonymous):

yeah. you dont see it as well?

OpenStudy (experimentx):

i on second derivative.

OpenStudy (experimentx):

let's take help form wolfram.

OpenStudy (experimentx):

http://www.wolframalpha.com/input/?i=d%5E9%2Fdx%5E9%28arctan%28x%5E%283%29%2F4%29%29 this is something terribly ugly ... i think we have to look for another method.

OpenStudy (anonymous):

Okay..

OpenStudy (dumbcow):

couldn't you just do this since you want to evaluate at x=0 http://www.wolframalpha.com/input/?i=d^9%2Fdx^9%28arctan%28x^%283%29%2F4%29%29+at+x%3D0

OpenStudy (dumbcow):

aren't you glad we have software now so you don't have to do these by hand :|

OpenStudy (experimentx):

isn't there another method ... it would take a week by hand.

OpenStudy (dumbcow):

probably but i don't know of it off top of my head

OpenStudy (asnaseer):

You could write arctan as a series expansion to get:\[\arctan(\frac{x^3}{4})=\frac{x^3}{4}-\frac{x^9}{192}+\frac{x^{15}}{5120}+...\]and then notice that the 9th derivative at x=0 will be zero for all terms except the term involving \(x^9\) which will end up as:\[-\frac{9!}{192}=-1890\]

OpenStudy (anonymous):

this seems correct I think.

OpenStudy (experimentx):

this was genius

OpenStudy (anonymous):

yep. thank you for your help!! experimentx

OpenStudy (experimentx):

Ah yes only up to third derivative http://www.wolframalpha.com/input/?i=d%5E3%2Fdx%5E3%28arctan%28x%5E%283%29%2F4%29%29+at+x%3D0